OpenStudy (anonymous):

what note does this end up to be.. Quarter note + sixteenth note +sixteenth note plus quarter note + eighth note?

OpenStudy (dmezzullo):

Close this and post it in the music group plz :)

OpenStudy (anonymous):

Wheres the music group?

OpenStudy (tkhunny):

4, 16, 16, 4, 8 Add the 16's 4, 8, 4, 8 Add the 8's 4, 4, 4 Add the 4's 3/4 or a dotted half. Math AND Music!

OpenStudy (austinl):

This is math @dmezzullo . \[\frac{1}{4} + \frac{1}{16} + \frac{1}{16} + \frac{1}{4} + \frac{1}{8}\]

OpenStudy (dmezzullo):

Well sorry

OpenStudy (austinl):

\[\frac{4}{16} + \frac{1}{16} + \frac{1}{16} + \frac{4}{16} + \frac{2}{16} = \frac{4 + 1 + 1 + 4 + 2}{16} = \frac{12}{16} = \frac{3}{4}\]

OpenStudy (anonymous):

@tkhunny Thanks..and what would that look like? A half note ? quarter note?...

OpenStudy (tkhunny):



OpenStudy (austinl):

The dotted half note.

OpenStudy (anonymous):

Thank you guys so muuchh ! it's correct

OpenStudy (anonymous):

@austinL @tkhunny

OpenStudy (austinl):

Very cool!

OpenStudy (tkhunny):

Well, yeah. You were expecting incorrect? :-)